OpenStudy (anonymous):

Find the solution to the equation 64^(4 – x) = 4^(2x)

OpenStudy (campbell_st):

rewrite 64 as a power of 4 \[(4^3)^{4 - x} = 4^{2x}\] this can be simplified using power of a power law \[4^{12 - 3x} = 4^{2x}\] since the bases are the same equate the powers 12 - 3x = 2x add 3x to both sides 12 = 5x just solve for x

OpenStudy (anonymous):

So x would be equal to 12/5

OpenStudy (campbell_st):

correct

OpenStudy (anonymous):

Thanks :D

OpenStudy (campbell_st):

hope it helps

OpenStudy (anonymous):

May you help me with finding the solution to the equation of log16 32 = x + 5.

OpenStudy (campbell_st):

start by looking at the left hand side and simplify the base 16 logs 32 = 2 x 16 so log 32 = log (2 x 16) = log2 + log 16 \[\log _{16} 16 = 1\] and \[\log _{16} 2 = \log _{16} \sqrt[4]{16} = \frac{1}{4}\] so the problem becomes \[\frac{1}{4} + 1 = x + 5\] or\[1\frac{1}{4} = x + 5\] solve for x